how long is a wire reaching from the top of a 20 foot pole to a point on the ground 14 feet from teh base of the pole?
Answer by checkley77(12569)

You have a right triangle with base=14 & hight=20
a^2+b^2=c^2
14^2+20^2=x^2
196+400=x^2
x^2=596
x=sqrt596
x=24.41 is the length of the wire.